Solve and show all work, please. Thanks!
A forester is 50 m from the base of a tree and measures the angle between the ground and the top of the tree (angel of elavation) to be 76 degrees. Find the height of the tree to the nearest meteer.

Answer by venugopalramana(3286)

A forester is 50 m from the base of a tree and measures the angle between the ground and the top of the tree (angel of elavation) to be 76 degrees. Find the height of the tree to the nearest meteer
LET F BE THE POSITION OF FORESTER ON GROUND AND GT BE THE TREE WITH G ON THE GROUND AND T ON TOP OF THE TREE.
FGT IS RIGHT ANGLED TRIANGLE WITH ANGLE FGT =90
FG=50 M.
ANGLE TFG=76
HT OF TREE =H =GT
HENCE
TAN (TFG)= GT/FG=H/50
H=50*TAN(TFG)=200 M.